A bus left a petrol station at 9.20 a.m. and travelled at an average speed of 75 km/h to a town N. At 9.40 a.m. a taxi travelling at an average speed of 95 km/h, left the same petrol station and followed the same route as the bus. Determine the distance, from the petrol station covered by the taxi at the time it caught up with the bus.
Answer by greenestamps(13200)



When the taxi leaves, the bus has been traveling for 20 minutes (1/3 of an hour) at 75km/h, traveling a distance of (1/3)(75) = 25km.

The rate at which the taxi catches up to the bus is the difference of their speeds -- 95-75 = 20km/hr.

The time required to make up the distance of 25km at a rate of 20km/h is 25/20 = 5/4 = 1.25 hours or 1 hour 15 minutes.

When the taxi catches up to the bus, it has been traveling 5/4 hours at a speed of 95km/h; the distance it has traveled is (5/4)(95) = 475/4 = 118.75km.

ANSWER: 118.75km